OpenStudy (anonymous):

Solve: 4 over 5 x + 4 over 3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

jigglypuff314 (jigglypuff314):

4/5 x + 4/3 = 2x multiply through by 15 *15 *15 *15 12x + 20 = 30x subtract 12x from both sides -12x -12x 20 = 18x divide both sides by 18 /18 /18 x = 20/18 = (can be simplified)